# Forecasting Treatment Response with Deep Pharmacokinetic Encoders

### v7

@@ Appendix A Supplemental Methods
Data and Code Availability
We use open-source simulated and real-world data that is available to other researchers. The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository [Xie, 2018, Visentin et al., 2016]. The real-world data includes the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020]. Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu [2020].
Implementations of models used in our work were obtained from the open-source Neuralforecast [Olivares et al., 2022] and StatsForecast [Garza et al., 2022] libraries. We implemented our method using a copy of the Neuralforecast repository to leverage their model training framework. Our research code, along with the Neuralforecast models used in our research, is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
@@ A.1 Data
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) [Xie, 2018, Visentin et al., 2016] provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes. Clinical parameters include information on age and weight, and biological parameters include information on insulin-glucose kinetics, such as absorption constants. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. In a similar approach to [Rubin-Falcone et al., 2022], the meal schedule used to generate simulated data was based on the Harrison-Benedict equation [Arthur Harris and Gano Benedict, 21919] with approximately 3 meals per day and no additional snacks. Height data was not provided in the patient parameters, and so was estimated at 140cm, 170cm, and 175cm for adults, adolescents, and children, respectively.
@@
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020]. Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The dataset consists of both female (n=5) and male (n=8) subjects with unspecified ages within the ranges of 20-40, 40-60, and 60-80. Information on insulin pump model and type of insulin is also provided.
@@
Fig. 7 shows the first 10 days of data for 6 subjects from the simulated dataset. Fig. 8 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset. Figs. 9 and 10 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively. Table 2 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
@@ A.2 Models
The baseline models used in this work have been extensively used in previous blood glucose forecasting work. For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies [Qian et al., 2021, Li et al., 2019a, Fox et al., 2018, Rabby et al., 2021, Rubin-Falcone et al., 2020, 2022, Miller et al., 2020]. Similarly, CNNs [Xie and Wang, 2020, Li et al., 2019a, b] and MLPs [McShinsky and Marshall, 2020, Miller et al., 2020] have been utilized in proposed models and as baselines. NHITS achieves state-of-the-art performance, surpassing RNN-based, transformer-based, and other deep learning time series decomposition models (such as N-BEATS [Oreshkin et al., 2020]) in long-horizon forecasting tasks [Challu et al., 2022].
@@
Exponential Smoothing (ETS) The ETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. [Hyndman, 2008].
@@
Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation [Nair and Hinton, 2010, Fukushima, 1975, Rosenblatt, 1958].
@@
Neural Hierarchical Interpolation for Time Series (NHITS) A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands. The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation [Challu et al., 2022].
@@
Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states. The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts.[Elman, 1990, Cho et al., 2014]
@@
Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.[Sak et al., 2014]
@@
Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts. To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions [Bai et al., 2018, van den Oord et al., 2016].
@@ A.3 Model Training and Hyperparameters
Models were trained using Adam optimizer [Kingma and Ba, 2017]. Optimal hyperparameters were selected based on a validation set with cross-validation. Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected. Huber estimator was selected as a robust loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)$ =$=\begin{cases}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta\cdot(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{cases}$ (8)
Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set. The deep learning models were trained using an NVIDIA A100 Tensor Core GPU.
For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 3. Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 3. We tuned the hyperparameters of our model using ‘HyperOptSearch’ with the Tree-structured Parzen Estimator (TPE) algorithm and 20 samples.
@@
We tune additional specific hyperparameters for each architecture. For the MLP  we tune the number of hidden units per layer ([128,256]); for the LSTM  the hidden size ([128,256]); for the TCN the hidden size ([128,256]) and number of dilations ([2,4,8]); for the NHITS we did not tune additional hyperparameters. The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
A.4 Metrics
We use the following metrics to evaluate model performance. Here, $H$ refers to the forecast horizon.
Mean Absolute Error (MAE)
$MAE(y_{t+1,t+H},\hat{y}_{t+1,t+H})=\frac{1}{H}\sum_{i=t+1}^{t+H}|y_{i}-\hat{y}_{i}|$
Root Mean Square Error (RMSE)
$RMSE(y_{t+1,t+H},\hat{y}_{t+1,t+H})=\sqrt{\frac{1}{H}\sum_{i=t+1}^{t+H}(y_{i}-\hat{y}_{i})^{2}}$
A.5 Propositions
@@ A.6 Algorithm
The algorithm initializes k as a network embeding weight using a value $\textbf{k}^{*}$ from the specified range $[0,1]$ at the beginning of each epoch. This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library [Liaw et al., 2018]. For each training step, the algorithm iterates over batches of data. It computes a context vector c based on the input data x, k, and other parameters. It then updates the hyperparameters k for each subject using stochastic gradient descent (SGD) based on the loss function $\mathcal{L}$ between the predicted $\hat{\textbf{y}}$ and actual y values. After a defined number of training steps (determined by a validation step parameter), the algorithm performs a validation step. It calculates the total loss $\ell(\textbf{k})_{\text{total}}$ over the validation set by summing the loss $\ell(\textbf{k})$ and the current batch loss $\mathcal{L}(\textbf{y},\hat{\textbf{y}})$. It then updates the hyperparameter vector $\textbf{k}^{*}$ to minimize the total loss. Finally, the algorithm returns the final k values.
@@
$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
@@
$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
@@ Appendix B Pharmacokinetic Background
From PK literature, exponential decay functions are commonly used to model drug concentration over time for intravenous injections, where the drug is directly injected into the bloodstream, leading to instantaneous effects [Rosenbaum, 2017, Angelika et al., 2005]. Alternatively, subcutaneous injections involve absorption and elimination processes where the peak onset time may not be instantaneous depending on the drug profile. Regarding knowledge on concentration-time profiles established through empirical and invasive studies, we know that concentration-time profiles for subcutaneous injections adhere to right-skewed curves. The degree of right-skew in the concentration curve will differ depending on the type of insulin, as illustrated in Fig. 2, along with various patient-specific factors.
@@
Insulin analogs are modified forms of insulin designed to exhibit certain PK properties, such as changes in drug onset and duration of action. Fig. 2 is adapted from PK literature and depicts PK modeling of plasma insulin concentration over time for different insulin analogs [Angelika et al., 2005]. Generally, bolus doses are quick-acting and, thus, often taken before meals. Contrastingly, basal insulin is longer-acting and serves to regulate blood glucose levels throughout the day. However, continuous glucose monitors (CGMs) and insulin pumps generally only hold and administer one type of insulin, such as rapid-acting insulin. For these devices, ‘bolus’ and ‘basal’ are terms used to refer to different timings of injections for different purposes. Bolus doses are administered discretely, generally around meals, while basal doses are administered continuously at a defined frequency, such as hourly, to regulate blood glucose levels throughout the day.
@@
In Fig. 2, the area under the concentration curve measures drug bioavailability, the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Bioavailability varies based on many factors, including type of medication, dosage, and patient-specific characteristics. For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent [Søeborg et al., 2012]. However, for subcutaneous injections, bioavailability is typically lower due to unintended drug absorption or elimination.
@@
Linear pharmacokinetics refers to a situation where the relationship between the dose of a drug and its concentration in the body is linear [Rosenbaum, 2017]. In linear pharmacokinetics, the area under the concentration curve is directly proportional to the dose of the drug. If you double the dose of a drug, the area under the curve will also double. Linearity results from the assumption that the drug processes exhibit first-order kinetics and that the PK parameters remain constant with dose, indicating no significant changes in the body’s ability to eliminate the drug with higher doses. The majority of drugs used in clinical practice are assumed to follow linear pharmacokinetics [Rosenbaum, 2017].
@@ C.1 Patient-specific (Local) Models
Table 4: MAE computed across forecast horizon values using local and global models.
ID Exogenous PK
@@
Figure 12: The NHITS-PK hybrid global-local model (Hybrid) trained on all OhioT1DM subject data (blue, solid) has a lower MAE, on average, than NHITS-PK models trained on individual subjects, or local models (purple, dashed) .
C.2 Forecast Comparison
Figure 13: Blood glucose forecasts of the final horizon point during 5 randomly selected hypoglycemic threshold periods (blood glucose $\geq$ 180) for 4 randomly selected OhioT1dm patients. The NHITS-PK hybrid global-local model (orange, dashed) produces improved forecasts of the ground truth (blue, solid) compared to the NHITS baseline model with sparse exogenous features (yellow, dash-dotted).
C.3 Forecasting Response to Treatment
Predicting treatment effects using accurate forecasting algorithms can assist individuals with diabetes and their care providers in making informed decisions about the ideal medication type and dosage, leading to improved glycemic control and reduced risk of both short- and long-term complications. Assessing a forecasting model’s ability to effectively capture the impact of variables on the target signal is essential for ensuring the reliability of forecasts. We examine whether the NHITS-PK model can accurately capture the treatment effects of insulin on blood glucose by generating forecasts during time periods for three scenarios: with the original bolus insulin doses, with bolus insulin doses removed, and with bolus insulin doses increased by a factor of 10. When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig. 14. The NHITS-PK model thus effectively learns the inhibitory treatment effects of insulin on blood glucose.
Figure 14: Treatment effect of insulin on the NHITS-PK hybrid global-local model. Each point in the plot corresponds to the final horizon point. Forecasts with original bolus doses (orange, solid), removing bolus insulin doses results in higher forecasts (blue, dashed), and augmenting doses results in lower blood glucose forecasts (yellow, dotted).
C.4 Forecasting with Dose-Independent Parameters
[…]
We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values. However, this approach led to worse performance on the OhioT1DM dataset. The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters [Xie, 2018, Visentin et al., 2016]. In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3. Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data. These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
Table 5: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)

### v8

@@ A.1 Data
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes. Clinical parameters include information on age and weight, and biological parameters include information on insulin-glucose kinetics, such as absorption constants. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. In a similar approach to (Rubin-Falcone et al., 2022), the meal schedule used to generate simulated data was based on the Harrison-Benedict equation (Arthur Harris and Gano Benedict, 21919) with approximately 3 meals per day and no additional snacks. Height data was not provided in the patient parameters, and so was estimated at 140cm, 170cm, and 175cm for adults, adolescents, and children, respectively.
@@
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The dataset consists of both female (n=5) and male (n=8) subjects with unspecified ages within the ranges of 20-40, 40-60, and 60-80. Information on insulin pump model and type of insulin is also provided.
@@
Fig. 10 shows the first 10 days of data for 6 subjects from the simulated dataset. Fig. 11 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset. Figs. 12 and 13 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively. Table 4 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
@@ A.2 Models
The baseline models used in this work have been extensively used in previous blood glucose forecasting work. For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020). Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020) have been used in proposed models and as baselines.
@@
Exponential Smoothing (AutoETS) The AutoETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. (Hyndman, 2008).
DLinear (DLinear) A Linear model that leverages linear layers to model the trend and seasonal components. A moving average filter is used to separate the time series into its trend and seasonal components (Zeng et al., 2023).
@@
Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
@@
Neural Hierarchical Interpolation for Time Series (NHITS) A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands. The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation (Challu et al., 2022).
Neural Basis Expansion Analysis with Exogenous (NBEATSx) An MLP-based deep learning model that decomposes the input signal into trend and seasonality components using polynomial and harmonic basis projections. The NBEATSx extension enables modeling of exogenous variables in addition to the univariate signal (Oreshkin et al., 2020; Olivares et al., 2022a).
@@
Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states. The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts (Elman, 1990; Cho et al., 2014).
@@
Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.(Sak et al., 2014)
@@
Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts. To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions (Bai et al., 2018; van den Oord et al., 2016).
Informer (Informer) (Zhou et al., 2021) A Transformer-based deep learning model with a multi-head attention mechanism that uses autoregressive features from a convolution network. A ProbSparse self-attention mechanism and a self-attention distilling process are leveraged to reduce computational complexity.
Patch time series Transformer (PatchTST) (Nie et al., 2023) A Transformer-based deep learning model with a multi-head attention mechanism. Input time series into sub-series level patches as input tokens.
Temporal Fusion Transformer (TFT) An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies. TFT also leverages variable selection networks as well as a series of gating layers to suppress unnecessary parts of the architecture (Lim et al., 2021).
@@ A.3 Model Training and Hyperparameters
The deep learning models were trained using an NVIDIA A100 Tensor Core GPU. For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 5. Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 5. We tuned the hyperparameters of our model using ‘HyperOptSearch’ with the Tree-structured Parzen Estimator (TPE) algorithm and 20 samples.
@@
We tune additional specific hyperparameters for each architecture. For the MLP model, we tune the number of hidden units per layer ([128, 256]). For the LSTM model, we tune the hidden size ([128, 256]). For the TCN model, we tune the hidden size ([128, 256]) and number of dilations ([2, 4, 8]). For the PatchTST model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and patch length ([8, 16, 32, 64, 96, 128]). For the Informer model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and number of decoder layers ([1, 2]). For the TFT model, we tune the hidden size ([128, 256, 512]) and number of attention heads ([4, 8, 16]). For the NBEATSx and NHITS models, we did not tune additional hyperparameters. The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
A.4 Propositions
@@ A.5 Algorithm
The algorithm initializes k as a network embeding weight using a value $\textbf{k}^{*}$ from the specified range $[0,1]$ at the beginning of each epoch. This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library (Liaw et al., 2018). For each training step, the algorithm iterates over batches of data. It computes a context vector c based on the input data x, k, and other parameters. It then updates the hyperparameters k for each subject using stochastic gradient descent (SGD) based on the loss function $\mathcal{L}$ between the predicted $\hat{\textbf{y}}$ and actual y values. After a defined number of training steps (determined by a validation step parameter), the algorithm performs a validation step. It calculates the total loss $\ell(\textbf{k})_{\text{total}}$ over the validation set by summing the loss $\ell(\textbf{k})$ and the current batch loss $\mathcal{L}(\textbf{y},\hat{\textbf{y}})$. It then updates the hyperparameter vector $\textbf{k}^{*}$ to minimize the total loss. Finally, the algorithm returns the final k values.
@@
$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
@@
$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
@@ Appendix B Pharmacokinetic Background
From PK literature, exponential decay functions are commonly used to model drug concentration over time for intravenous injections, where the drug is directly injected into the bloodstream, leading to instantaneous effects (Rosenbaum, 2017; Angelika et al., 2005). Alternatively, subcutaneous injections involve absorption and elimination processes where the peak onset time may not be instantaneous depending on the drug profile. Regarding knowledge on concentration-time profiles established through empirical and invasive studies, we know that concentration-time profiles for subcutaneous injections adhere to right-skewed curves. The degree of right-skew in the concentration curve will differ depending on the type of insulin, as illustrated in Fig. 2, along with various patient-specific factors.
@@
Insulin analogs are modified forms of insulin designed to exhibit certain PK properties, such as changes in drug onset and duration of action. Fig. 2 is adapted from PK literature and depicts PK modeling of plasma insulin concentration over time for different insulin analogs (Angelika et al., 2005). Generally, bolus doses are quick-acting and, thus, often taken before meals. Contrastingly, basal insulin is longer-acting and serves to regulate blood glucose levels throughout the day. However, continuous glucose monitors (CGMs) and insulin pumps generally only hold and administer one type of insulin, such as rapid-acting insulin. For these devices, ‘bolus’ and ‘basal’ are terms used to refer to different timings of injections for different purposes. Bolus doses are administered discretely, generally around meals, while basal doses are administered continuously at a defined frequency, such as hourly, to regulate blood glucose levels throughout the day.
@@
In Fig. 2, the area under the concentration curve measures drug bioavailability, the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Bioavailability varies based on many factors, including type of medication, dosage, and patient-specific characteristics. For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent (Søeborg et al., 2012). However, for subcutaneous injections, bioavailability is typically lower due to unintended drug absorption or elimination.
@@
Linear pharmacokinetics refers to a situation where the relationship between the dose of a drug and its concentration in the body is linear (Rosenbaum, 2017). In linear pharmacokinetics, the area under the concentration curve is directly proportional to the dose of the drug. If you double the dose of a drug, the area under the curve will also double. Linearity results from the assumption that the drug processes exhibit first-order kinetics and that the PK parameters remain constant with dose, indicating no significant changes in the body’s ability to eliminate the drug with higher doses. The majority of drugs used in clinical practice are assumed to follow linear pharmacokinetics (Rosenbaum, 2017).
@@ C.1 Patient-specific (Local) Models
Table 6: MAE computed across forecast horizon values using local and global models for the OhioT1DM dataset. Best results are in bold.
ID w/ Exog. (Sparse) w/ Exog. (PK)
@@
C.2 Forecasting with Dose-Independent Parameters
Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180). Result for models with (w/) and without (w/o) exogenous variables are shown. Best results are in bold.